How many solutions does -14x+7y=21
-8x+4y=12 have?

Answer by ewatrrr(24785)

-8x+4y=12 0r y = 2x + 3
-14x+7y=21
-14x + 7(2x+3) = 21
21 = 21 Infinite Solutions...same line